# Silenced Biases: The Dark Side LLMs Learned to Refuse

### v1

@@ 6 Discussion
The findings of this paper suggest that safety-alignment induced refusals for silenced biases do not outright remove them. We uncover that these biases remain latent within model activations and can be revealed quite easily. This is important since these biases might affect normal day-to-day interactions with the LLM. An LLM can potentially identify a user’s demographic affiliation, and start responding with biased intent, without the user ever knowing.
To the best of our knowledge, this is the first study to empirically demonstrate that refusal behavior conceals bias, offering a framework to audit silenced biases. By exposing this blind spot, we open a new direction for fairness and bias evaluation, one that accounts for alignment-induced silence. The broader impact of this work lies in enabling better debiasing strategies, deeper understanding of alignment’s limitations, and a more complete view of model behavior. Future work should focus on building tools that go beyond surface-level outputs to uncover the biased associations still present in the model’s latent space. It is equally important to better understand alignment training itself, which often makes models appear unbiased while still preserving problematic stereotypes internally. We discuss limitations and ethical considerations in Appendices B and A.
@@ C.2 Figure 2 - Compliance and Refusal Shift
Figure 1 is heavily inspired by Levi et al. (2025). To generate it, we first sampled both harmful and harmless prompts from the training split of Arditi (2024) and computed a refusal direction using the method described in the main text. We then sampled harmful and harmless prompts from the test split of the same dataset and extracted their activation vectors, taking the representation at the final layer and final token position.
@@ D.1 Benchmark Direction Similarity to Refusal Direction
In this evaluation, we aim to quantify the harmfulness of various benchmarks by measuring the extent to which they elicit refusals from the model. To do so, we leverage the refusal direction and compute similarity to it, following the methodology of Arditi (2024). Specifically, we calculate the refusal direction using the harmful and harmless training sets, sampling 128 prompts from each, as described in the original work and later done in our experiments.
To assess the similarity of a given benchmark to the refusal direction, we first sample 128 prompts from the benchmark. For each prompt, we compute its activation similarity to the refusal direction across all layers, and then average these similarities across all prompts. For the harmful and harmless sets, we use the test set provided by Arditi (2024).
@@ D.3 Differing Representation of Biased and Unbiased Responses
In Figure 8, we present additional visualizations of separations between biased and unbiased classes, emphasizing this is not a specific category not model layer occurrence.
@@ F.2 Experimental results
In Figures 12 and 13 we provide the full measures across all demographic categories in SBB, including DPD and KL divergence.
@@
Figures 14 and 15 shows the average Jaccard similarity over all pairs of directions (numbered 0 through 9) for each evaluated LLM. The consistently high similarity values indicate that the different directions mainly implement refusal steering without introducing substantial new bias in the content of responses. Residual deviations from perfect agreement (the minimum observed similarity across all models is 0.75) can be attributed largely to sampling variability in the generation process.

### v2

@@ 6 Discussion
The findings of this paper suggest that safety-alignment induced refusals for silenced biases do not outright remove them. We uncover that these biases remain latent within model activations and can be revealed quite easily. This is important since these biases might affect normal day-to-day interactions with the LLM. An LLM can potentially identify a user’s demographic affiliation and start responding with biased intent, without the user ever knowing.
To the best of our knowledge, this is the first study to empirically demonstrate that refusal behavior conceals bias, offering a framework to audit silenced biases. By exposing this blind spot, we open a new direction for fairness and bias evaluation, one that accounts for alignment-induced silence. We encourage the extension of SBB to more complex scenarios than QA, such as reasoning chains, long-form generation, or agentic behavior. The broader impact of this work lies in enabling better debiasing strategies, a deeper understanding of alignment’s limitations, and a more complete view of model behavior. Future work should focus on building tools that go beyond surface-level outputs to uncover the biased associations still present in the model’s latent space. It is equally important to better understand alignment training itself, which often makes models appear unbiased while still preserving problematic stereotypes internally. We discuss limitations and ethical considerations in Appendices B and A.
Acknowledgments
This research was partially supported by the United States-Israel Binational Science Foundation (Grant No. 2024101) and by the Israel Science Foundation (Grant No. 934/25).
@@ C.2 Figure 2 - Compliance and Refusal Shift
Figure 1 is heavily inspired by Levi et al. (2025). To generate it, we first sampled both harmful and harmless prompts from the training split of Arditi et al. (2024) and computed a refusal direction using the method described in the main text. We then sampled harmful and harmless prompts from the test split of the same dataset and extracted their activation vectors, taking the representation at the final layer and final token position.
@@ D.1 Benchmark Direction Similarity to Refusal Direction
In this evaluation, we aim to quantify the harmfulness of various benchmarks by measuring the extent to which they elicit refusals from the model. To do so, we leverage the refusal direction and compute similarity to it, following the methodology of Arditi et al. (2024). Specifically, we calculate the refusal direction using the harmful and harmless training sets, sampling 128 prompts from each, as described in the original work and later done in our experiments.
To assess the similarity of a given benchmark to the refusal direction, we first sample 128 prompts from the benchmark. For each prompt, we compute its activation similarity to the refusal direction across all layers, and then average these similarities across all prompts. For the harmful and harmless sets, we use the test set provided by Arditi et al. (2024).
@@ D.3 Differing Representation of Biased and Unbiased Responses
In Figure 8, we present additional visualizations of separations between biased and unbiased classes, emphasizing this is not a specific category not model layer occurrence. Moreover, in Figure 9, we present the silhouette coefficient on these biased and unbiased clusters on the original activations across layers and subjects. We notice the activations in the first layer are the most large and separate. More experiments should be conducted, in order to explore the layer relationship of biases across more LLMs.
@@
Figure 9: Silhouette scores of biased and unbiased clusters, per subject and layer, on Llama2-7B.
@@ F.2 Experimental results
In Figures 13 and 14 we provide the full measures across all demographic categories in SBB, including DPD and KL divergence.
@@
Figures 15 and 16 shows the average Jaccard similarity over all pairs of directions (numbered 0 through 9) for each evaluated LLM. The consistently high similarity values indicate that the different directions mainly implement refusal steering without introducing substantial new bias in the content of responses. Residual deviations from perfect agreement (the minimum observed similarity across all models is 0.75) can be attributed largely to sampling variability in the generation process.